# Statistical Inference of Constrained Stochastic Optimization via Sketched Sequential Quadratic Programming

### v4

@@ 1 Introduction
where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a stochastic objective function, $F(\cdot;\xi):\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a realization with a random variable $\xi\sim{\mathcal{P}}$, and $c:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ provides deterministic equality constraints. Problems of this form appear widely in a variety of applications in statistics and machine learning, including constrained $M$-estimation (Geyer1991Constrained; Geyer1994Asymptotics; Wets1999Statistical), multi-stage stochastic optimization (Dantzig1993Multi; Veliz2014Stochastic), physics-informed neural networks (Karniadakis2021Physics; Cuomo2022Scientific), and algorithmic fairness (Zafar2019Fairness). In practice, the random variable $\xi$ corresponds to a data sample; $F({\bm{x}};\xi)$ is the loss occurred at sample $\xi$ when using parameter ${\bm{x}}$ to fit the model; and $f({\bm{x}})$ is the expected loss. Deterministic constraints are prevalent in real examples, which can encode prior model information, address identifiability issue, and/or reduce searching complexity.
@@
where $\mathcal{L}({\bm{x}},{\bm{\lambda}})=f({\bm{x}})+{\bm{\lambda}}^{T}c({\bm{x}})$ is the Lagrangian function with ${\bm{\lambda}}\in\mathbb{R}^{m}$ being the dual variables associated with the constraints, $\nabla_{{\bm{x}}}^{2}\mathcal{L}^{\star}$ is the Lagrangian Hessian with respect to ${\bm{x}}$ evaluated at $({\bm{x}}^{\star},{\bm{\lambda}}^{\star})$, and $G^{\star}=\nabla c({\bm{x}}^{\star})\in\mathbb{R}^{m\times d}$ is the constraints Jacobian. See (Shapiro2014Lectures, Chapter 5) for the result of (1.3) and Duchi2021Asymptotic; Davis2023Asymptotic for showing (1.3) attains the minimax optimality. Numerous methods can be applied to solve constrained ERM (1.2), including penalty methods, augmented Lagrangian methods, and sequential quadratic programming (SQP) methods (Nocedal2006Numerical).
@@
Given the prevalence of streaming datasets in modern problems, offline methods that require dealing with a large batch set in each step are less attractive. It is desirable to design fully online methods, where only a single sample is used in each step, and to perform online statistical inference by leveraging those methods. Without constraints, one can apply stochastic gradient descent (SGD) and its many variates, whose statistical properties (e.g., asymptotic normality) have been comprehensively studied from different aspects (Robbins1951stochastic; Kiefer1952Stochastic; Polyak1992Acceleration; Ruppert1988Efficient). However, unlike solving unconstrained stochastic programs, there are limited methods proposed for constrained stochastic programs (1.1) that enable online statistical inference. We refer to Section 2.2 for a detailed literature review. One potential exception is the projection-based SGD recently studied in Duchi2021Asymptotic; Davis2023Asymptotic. Although the literature has shown that projected methods also exhibit asymptotic normality, there are two major concerns when applying these methods for practical statistical inference. First, it is unclear how to online estimate the limiting covariance matrix based on the projected iterates, which is highly nontrivial as explored in vanilla SGD method (Chen2020Statistical; Zhu2021Online). Second, there are prevalent scenarios where the projection operator becomes intractable. For example, when the constraint function $c({\bm{x}})$ is nonlinear and nonconvex, as in physics-informed neural networks (cf. Section 2.1), the projection onto the manifold $\{{\bm{x}}\in\mathbb{R}^{d}:c({\bm{x}})={\bm{0}}\}$ is generally intractable. Additionally, if we only have local information about the constraint function $c({\bm{x}})$ (e.g., function evaluation and Jacobian) at any given point ${\bm{x}}$, as in CUTEst benchmark nonlinear problems (cf. Section LABEL:sec:5), the projection operator is not computable either, which requires a global characterization of the constraint set.
@@
To perform online statistical inference of Problem (1.1) without relying on projections, we draw inspiration from a recent growing series of literature in numerical optimization, which develops various stochastic sequential quadratic programming (StoSQP) methods for (1.1). The SQP methods can be regarded as second-order Newton’s method applied to the Karush-Kuhn-Tucker (KKT) conditions. In particular, the StoSQP method computes a stochastic Newton direction in each iteration by solving a quadratic program, whose objective model is estimated using the new sample. Then, the method selects a proper stepsize to achieve a sufficient reduction on the merit function, which balances the optimality and feasibility of the iterates. We refer to Na2022adaptive; Na2023Inequality; Berahas2021Sequential; Berahas2023Stochastic; Fang2024Fully and references therein for recent StoSQP designs and their promising performance on various problems. The aforementioned literature established the global convergence of StoSQP methods, where the KKT residual $\|\nabla\mathcal{L}({\bm{x}}_{t},{\bm{\lambda}}_{t})\|$ converges to zero almost surely or in expectation from any initialization. However, in contrast to Duchi2021Asymptotic; Davis2023Asymptotic studying projection-based methods, these literature overlooked the statistical properties and failed to quantify the uncertainty inherent in StoSQP methods, which is yet crucial for applying the methods on online inference tasks. Thus, we pose the following question:
@@
In this paper, we answer this question by complementing the global convergence guarantees and establishing the local asymptotic properties of existing StoSQP methods. Specifically, we focus on an Adaptive Inexact StoSQP scheme, referred to as AI-StoSQP. By adaptive we mean that the scheme inherits the critical merit of numerical StoSQP designs (Berahas2021Sequential; Curtis2021Inexact; Berahas2023Stochastic), allowing for an adaptive stepsize ${\bar{\alpha}}_{t}$ for the Newton direction. In other words, we do not compromise the adaptivity of StoSQP to establish the local convergence guarantees. By inexact we mean that the scheme further reduces the computational cost of StoSQP methods by employing an iterative sketching solver to inexactly solve the Newton system in each step (Strohmer2008Randomized; Gower2015Randomized; Pilanci2017Newton). Solving Newton systems is considered the most computationally expensive step of second-order methods; and randomized solvers offer advantages over deterministic solvers by requiring less flops and memory when equipped with proper sketching matrices (e.g., sparse sketches). Notably, we perform a constant number of sketching steps; thus, the per-iteration computational cost remains fixed even near stationarity.
@@
This implies that (i) if we solve Newton systems exactly, then online StoSQP estimator (even with adaptive stepsizes) achieves the same estimation efficiency as offline $M$-estimator. In fact, if we focus solely on the primal variables ${\bm{x}}$, the marginal covariance of ${\Xi^{\star}}={\Omega^{\star}}$ also matches the limiting covariance of online projection-based estimators established in Duchi2021Asymptotic; Davis2023Asymptotic, which is known to be asymptotic minimax optimal. (ii) If we solve Newton systems inexactly, then the sketching solver hurts the asymptotic optimality of StoSQP as ${\Xi^{\star}}\succeq{\Omega^{\star}}$. Fortunately, the hurt is tolerable as indicated by the bound (cf. Corollary LABEL:cor:3)
@@
Structure of the paper. We introduce some motivating examples of Problem (1.1) and provide a literature review in Section 2. Then, we introduce AI-StoSQP in Section 3, and prove the global almost sure convergence with iteration complexity in Section 4. Asymptotic normality with covariance estimation is established in Section LABEL:sec:4. Experiments and conclusions are presented in Sections LABEL:sec:5 and LABEL:sec:6, respectively. We defer all the proofs to the appendices.
@@ 2.1 Motivating examples
Let $\xi_{t}=(\xi_{{\bm{a}}_{t}},\xi_{b_{t}})$ be the $t$-th sample, where $\xi_{{\bm{a}}_{t}}\in\mathbb{R}^{d}$ is the feature vector independently drawn from some multivariate distribution, and $\xi_{b_{t}}$ is the response. We consider different models, such as
@@
Then, we can verify that ${\bm{x}}^{\star}=\mathop{\mathrm{argmin}}_{{\bm{x}}}f({\bm{x}})=\mathop{\mathrm{argmin}}_{{\bm{x}}}\mathbb{E}[F({\bm{x}};\xi)]$. In many cases, we access prior information about the model parameters, which is encoded as constraints. For example, in portfolio selection, ${\bm{x}}$ represents the portfolio allocation vector that satisfies ${\bm{x}}^{T}{\bm{1}}=1$ (total allocation is 100%). We may also fix the target percentage on each sector or each region, which is translated into the constraint $A{\bm{x}}={\bm{d}}$. See (Fan2007Variable, (4.3), (4.4)), (Fan2012Vast, (2.1)), (Du2022High, (1)) and references therein for such applications. In principle component analysis and semiparametric single/multiple index regressions, we enforce ${\bm{x}}$ to have a unit norm to address the identifiability issue, leading to a nonlinear constraint $\|{\bm{x}}\|^{2}=1$. We point to Kirkegaard1972POSITRONFIT; Kaufman1978Method; Sen1979Asymptotic; Nagaraj1991Estimation; Dupacova1988Asymptotic; Shapiro2000asymptotics; Na2019High; Na2021High for various examples of linearly/nonlinearly constrained estimation problems.
@@
Here, we abuse the notation ${\bm{0}}$ to denote either a zero mapping of $\mathcal{C}^{\infty}(\mathcal{A},\mathcal{B})$ or a zero element of $\mathcal{B}$. For more details on this problem formulation, see (Lu2021Physics, (2.3))(Krishnapriyan2021Characterizing, (2)), and references therein.
@@ 2.2 Related literature and contribution
On the other hand, a growing body of literature leverages optimization procedures to facilitate online inference, starting with Robbins1951stochastic; Kiefer1952Stochastic and continuing through Robbins1971convergence; Fabian1973Asymptotically; Ermoliev1983Stochastic. To study the asymptotic distribution of stochastic gradient descent (SGD), Ruppert1988Efficient and Polyak1992Acceleration averaged SGD iterates and established the optimal central limit theorem rate. Toulis2017Asymptotic designed an implicit SGD method and showed the asymptotics of averaged implicit SGD iterates. Li2018Statistical designed an inference procedure for constant-stepsize SGD by averaging the iterates with recurrent burn-in periods. Mou2020linear further showed the asymptotic covariance of constant-stepsize SGD with Ployak-Ruppert averaging. Liang2019Statistical designed a moment-adjusted SGD method and provided non-asymptotic results that characterize the statistical distribution as the batch size of each step tends to infinity. Chen2020Statistical and Zhu2021Online proposed different covariance matrix estimators constructed by grouping SGD iterates. Additionally, Chen2021First designed a distributed method for the inference of non-differentiable convex problemsRoy2023Online analyzed a batch-mean covariance estimator under a Markovian sampling setup; and Duchi2021Asymptotic and Davis2023Asymptotic applied projection-based SGD methods for the inference of inequality-constrained convex problems. The aforementioned literature all studied first-order methods with deterministic stepsizes.
@@ 3 Adaptive Inexact StoSQP Method
$\bullet$ Step 1: Estimate the gradient and Hessian. We realize a sample $\xi_{t}\sim{\mathcal{P}}$ and estimate
@@
We note that $B_{t}$ and $\Delta_{t}$ are deterministic given $({\bm{x}}_{t},{\bm{\lambda}}_{t})$, meaning they do not depend on the sample $\xi_{t}$. Additionally, $\Delta_{t}$ can simply be Levenberg-Marquardt type regularization of the form $\delta_{t}I$ with suitably large $\delta_{t}>0$ (Nocedal2006Numerical). The Hessian regularization is standard for nonlinear problems, together with linear independence constraint qualification (LICQ, Assumption 4.1), ensuring that the quadratic program (3.1) is solvable. For convex problems, we just set $\Delta_{t}={\bm{0}}$, $\forall t$. See Bertsekas1982Constrained; Nocedal2006Numerical for various regularization approaches. FurthermoreNa2022Hessian showed that Hessian averaging accelerates the local rate of Newton’s method on unconstrained deterministic problems.
@@
Instead of solving the QP (3.1) exactly, we solve it inexactly by an iterative sketching solver. This approach proves more efficient than deterministic solvers, especially when equipped with suitable sketching matrices (Strohmer2008Randomized; Gower2015Randomized; Pilanci2017Newton). In particular, we generate a random sketching matrix $S\in\mathbb{R}^{(d+m)\times q}$, whose column dimension $q\geq 1$ can also be random, and transform the original large-scale linear system to the sketched, small-scale system as
@@
where $(\cdot)^{\dagger}$ denotes the Moore–Penrose pseudoinverse. One can set $q=1$ (i.e., using sketching vectors) so that $S_{t,j}^{T}K_{t}^{2}S_{t,j}$ reduces to a scalar and the pseudoinverse reduces to the reciprocal.
@@
A significant difference between randomized solvers and deterministic solvers is that the approximation error $\|{\bm{z}}_{t,j}-{\widetilde{\bm{z}}}_{t}\|$ of randomized solvers may not be monotonically decreasing as $j$ increases. This subtlety challenges both inference and convergence analysis. In classical optimization world, it is unanimously agreed that if the search direction is asymptotically close to the exact Newton direction (here ${\widetilde{\bm{z}}}_{t}$), then the algorithm will locally behave just like Newton’s method with a similar convergence rate. A precise characterization is called the Dennis-Moré condition (Dennis1974characterization). In our study, ${\bm{z}}_{t,\tau}$ can be far from ${\widetilde{\bm{z}}}_{t}$ for any $t$ and any large $\tau$.
@@
where $\{\beta_{t},\eta_{t}\}$ are upper and lower bound sequences and $\chi_{t}$ is the adaptivity gap. We do not require specific stepsize selection schemes beyond the condition (3.6) to achieve our online inference goals. The schemes proposed in Berahas2021Sequential; Berahas2023Stochastic; Curtis2021Inexact all adhere to the condition (3.6). We refer to (Berahas2021Sequential, Lemma 3.6) and (Curtis2021Inexact, (25),(28)) for details. Their numerical experiments suggest that adaptive random stepsizes offer promising empirical benefits over non-adaptive deterministic stepsizes (i.e., $\chi_{t}=0$). For sake of completeness, we present a selection scheme from Berahas2021Sequential in Appendix LABEL:app:1.

### v5

@@ 1 Introduction
[…]
@@
where $\mathcal{L}({\bm{x}},{\bm{\lambda}})=f({\bm{x}})+{\bm{\lambda}}^{T}c({\bm{x}})$ is the Lagrangian function with ${\bm{\lambda}}\in\mathbb{R}^{m}$ being the dual variables associated with the constraints, $\nabla_{{\bm{x}}}^{2}\mathcal{L}^{\star}$ is the Lagrangian Hessian with respect to ${\bm{x}}$ evaluated at $({\bm{x}}^{\star},{\bm{\lambda}}^{\star})$, and $G^{\star}=\nabla c({\bm{x}}^{\star})\in\mathbb{R}^{m\times d}$ is the constraints Jacobian. See (Shapiro2014Lectures, Chapter 5) for the result of (1.3), and Duchi2021Asymptotic and Davis2024Asymptotic for showing (1.3) attains the minimax optimality. Numerous methods can be applied to solve constrained ERM (1.2), including (exact) penalty methods, augmented Lagrangian methods, and sequential quadratic programming (SQP) methods (Nocedal2006Numerical).
@@
Given the prevalence of streaming datasets in modern problems, offline methods that require dealing with a large batch set in each step are less attractive. It is desirable to design fully online methods, where only a single sample is used in each step, and to perform online statistical inference by leveraging those methods. Without constraints, one can apply stochastic gradient descent (SGD) and its many variates, whose statistical properties (e.g., asymptotic normality) have been comprehensively studied from different aspects (Robbins1951stochastic; Kiefer1952Stochastic; Polyak1992Acceleration; Ruppert1988Efficient). However, unlike solving unconstrained stochastic programs, there are limited methods proposed for constrained stochastic programs (1.1) that enable online statistical inference. We refer to Section 2.2 for a detailed literature review. One potential exception is the projection-based SGD recently studied in Duchi2021Asymptotic; Davis2024Asymptotic. Although the literature has shown that projected methods also exhibit asymptotic normality, there are two major concerns when applying these methods for practical statistical inference.
  1. (a)
    It is unclear how to online estimate the limiting covariance matrix based on the projected primal iterates. Due to the absence of dual update, the joint primal-dual normality as in (1.3) is not (at least, immediately) achievable for projected methods. Even for the primal normality, the covariance matrix still depends on the dual solution through the Lagrangian Hessian (cf. (1.3)). However, due to intrinsic objective noise, simply using primal iterates and optimality conditions to solve for the dual solution does not yield a consistent dual estimator for the underlying plug-in covariance estimation. One possible resolution is to draw inspiration from long-run variance estimations of stationary processes, and design batch-means covariance estimators that utilize only the projected primal iterates themselves. That said, this approach is highly nontrivial for projected methods (because of the non-stationarity), as studied in the context of vanilla SGD methods (Chen2020Statistical; Zhu2021Online).
  2. (b)
    There are prevalent scenarios where the projection operator becomes intractable. For example, when the constraint function $c({\bm{x}})$ is nonlinear and nonconvex, as in physics-informed neural networks (cf. Section 2.1), the projection onto the manifold $\{{\bm{x}}\in\mathbb{R}^{d}:c({\bm{x}})={\bm{0}}\}$ is generally intractable. Additionally, if we only have local information about the constraint function $c({\bm{x}})$ (e.g., function evaluation and Jacobian) at any given point ${\bm{x}}$, as in CUTEst benchmark nonlinear problems (cf. Section LABEL:sec:5), the projection operator is not computable either, which requires a global characterization of the constraint set.
@@
To perform online statistical inference of Problem (1.1) without relying on projections, we draw inspiration from a recent growing series of literature in numerical optimization, which develops various stochastic sequential quadratic programming (StoSQP) methods for (1.1). The SQP methods can be regarded as second-order Newton’s methods applied to the Karush-Kuhn-Tucker (KKT) conditions. In particular, the StoSQP methods compute a stochastic Newton direction in each iteration by solving a quadratic program, whose objective model is estimated using the new sample. Then, the methods select a proper stepsize to achieve a sufficient reduction on the merit function, which balances the optimality and feasibility of the iterates. We refer to Na2022adaptive; Na2023Inequality; Berahas2021Sequential; Berahas2023Stochastic; Fang2024Fully and references therein for recent StoSQP designs and their promising performance on various problems. The aforementioned literature established the global convergence of StoSQP methods, where the KKT residual $\|\nabla\mathcal{L}({\bm{x}}_{t},{\bm{\lambda}}_{t})\|$ converges to zero almost surely or in expectation for any initialization. However, in contrast to Duchi2021Asymptotic; Davis2024Asymptotic studying projection-based methods, these literature overlooked the statistical properties and failed to quantify the uncertainty inherent in the StoSQP methods, which is yet crucial for applying these methods on online statistical inference tasks. Thus, we pose the following question:
@@
In this paper, we answer this question by complementing the global convergence guarantees and establishing the local asymptotic properties of existing StoSQP methods. Specifically, we focus on an Adaptive Inexact StoSQP scheme, referred to as AI-StoSQP. By adaptive we mean that the scheme inherits the critical merit of numerical StoSQP designs (Berahas2021Sequential; Curtis2021Inexact; Berahas2023Stochastic), allowing for an adaptive stepsize ${\bar{\alpha}}_{t}$ for the Newton direction. In other words, we do not compromise the adaptivity of StoSQP to establish the local convergence guarantees. By inexact we mean that the scheme further reduces the computational cost of StoSQP methods by applying an iterative sketching solver to inexactly solve the Newton system in each step (Strohmer2008Randomized; Gower2015Randomized; Pilanci2016Iterative; Pilanci2017Newton; Lacotte2020Optimal). Solving Newton systems is considered the most computationally expensive step of second-order methods; and randomized solvers offer advantages over deterministic solvers by requiring less flops and memory when equipped with proper sketching matrices (e.g., sparse sketches). Notably, we perform a constant number of sketching steps; thus, the per-iteration computational cost remains fixed even near stationarity.
@@
This implies that (i) if we solve Newton systems exactly, then online StoSQP estimator (even with adaptive stepsizes) achieves the same estimation efficiency as offline $M$-estimator. In fact, if we focus solely on the primal variables ${\bm{x}}$, the marginal covariance of ${\Xi^{\star}}\hskip-1.0pt=\hskip-1.0pt{\Omega^{\star}}$ also matches the limiting covariance of online projection-based estimators established in Duchi2021Asymptotic; Davis2024Asymptotic, which is known to be asymptotic minimax optimal (see Remark LABEL:rem:2). (ii) If we solve Newton systems inexactly, then the sketching solver hurts the asymptotic optimality of StoSQP as ${\Xi^{\star}}\succeq{\Omega^{\star}}$. Fortunately, the hurt is tolerable as seen from the bound (cf. Corollary LABEL:cor:3)
@@
Structure of the paper. We introduce some motivating examples of Problem (1.1) and provide a literature review in Section 2. Then, we introduce AI-StoSQP in Section 3 and prove the global almost sure convergence with iteration complexity in Section 4. Asymptotic normality with covariance estimation is established in Section LABEL:sec:4. Experiments and conclusions are presented in Sections LABEL:sec:5 and LABEL:sec:6, respectively. We defer all the proofs to the appendices.
@@ 2.1 Motivating examples
Let $\xi_{t}=(\xi_{{\bm{a}}_{t}},\xi_{b_{t}})$ be the $t$-th sample, where $\xi_{{\bm{a}}_{t}}\in\mathbb{R}^{d}$ is the feature vector independently drawn from some multivariate distribution and $\xi_{b_{t}}$ is the response. We consider different regression models, such as
@@
Then, we can verify that ${\bm{x}}^{\star}=\mathop{\mathrm{argmin}}_{{\bm{x}}}f({\bm{x}})=\mathop{\mathrm{argmin}}_{{\bm{x}}}\mathbb{E}[F({\bm{x}};\xi)]$. In many cases, we access prior information about the model parameters, which is encoded as constraints. For example, in portfolio selection, ${\bm{x}}$ represents the portfolio allocation vector that satisfies ${\bm{x}}^{T}{\bm{1}}=1$ (total allocation is 100%). We may also fix the target percentage on each sector or each region, which is translated into the constraint $A{\bm{x}}={\bm{d}}$. See (Fan2007Variable, (4.3, 4.4)), (Fan2012Vast, (2.1)), (Du2022High, (1)) and references therein for such applications. In principle component analysis and semiparametric single/multiple index regressions, we enforce ${\bm{x}}$ to have a unit norm to address the identifiability issue, leading to a nonlinear constraint $\|{\bm{x}}\|^{2}=1$. We point to Kaufman1978Method; Kirkegaard1972POSITRONFIT; Sen1979Asymptotic; Nagaraj1991Estimation; Dupacova1988Asymptotic; Shapiro2000asymptotics; Na2019High; Na2021High for various examples of linearly/nonlinearly constrained estimation problems. For some constrained estimation problems, projecting into the feasible set can be intractable. For instance, in factor analysis, researchers may estimate a covariance matrix $\Sigma$ under so-called tetrad constraints: $\Sigma_{i_{1}i_{2}}\Sigma_{i_{3}i_{4}}-\Sigma_{i_{1}i_{4}}\Sigma_{i_{2}i_{3}}=0$ for every set of four distinct variables $\{i_{1},i_{2},i_{3},i_{4}\}$ (Bollen2000tetrad; Drton2016Wald; Sturma2024Testing). For such highly nonlinear constraints, the linear-quadratic approximation performed in SQP can be a promising resolution.
@@
Here, we abuse the notation ${\bm{0}}$ to denote either a zero mapping of $\mathcal{C}^{\infty}(\mathcal{A},\mathcal{B})$ or a zero element of $\mathcal{B}$. For more details on this problem formulation, see (Lu2021Physics, (2.3))(Krishnapriyan2021Characterizing, (2)), and references therein. Due to the nonlinearity nature of NNs, projection-free methods are desired, and SQP can achieve competitive performance compared to penalty methods and augmented Lagrangian methods (Cheng2024Physics).
@@ 2.2 Related literature and contribution
On the other hand, a growing body of literature leverages optimization procedures to facilitate online inference, starting with Robbins1951stochastic; Kiefer1952Stochastic and continuing through Robbins1971convergence; Fabian1973Asymptotically; Ermoliev1983Stochastic. To study the asymptotic distribution of stochastic gradient descent (SGD), Ruppert1988Efficient and Polyak1992Acceleration averaged SGD iterates and established the optimal central limit theorem rate. Toulis2017Asymptotic designed an implicit SGD method and showed the asymptotics of averaged implicit SGD iterates. Li2018Statistical designed an inference procedure for constant-stepsize SGD by averaging the iterates with recurrent burn-in periods. Mou2020linear further showed the asymptotic covariance of constant-stepsize SGD with Ployak-Ruppert averaging. Liang2019Statistical designed a moment-adjusted SGD method and provided non-asymptotic results that characterize the statistical distribution as the batch size of each step tends to infinity. Chen2020Statistical and Zhu2021Online proposed different covariance matrix estimators constructed by grouping SGD iterates. Additionally, Chen2021First designed a distributed method for the inference of non-differentiable convex problems; Roy2023Online analyzed a batch-mean covariance estimator under a Markovian sampling setup; and Duchi2021Asymptotic and Davis2024Asymptotic applied projection-based SGD methods for the inference of inequality-constrained convex problems. The aforementioned literature all studied first-order methods with deterministic stepsizes.
@@
We briefly review the literature on sketched Newton methods. Compared to the works below, the sketching step in StoSQP is only a subroutine for solving linear-quadratic programs; a complete method also involves merit function reduction and stepsize selection (here, stepsize refers to that of StoSQP rather than the sketching solver). This paper focuses on uncertainty quantification and statistical inference of online (sketched) Newton methods, which differs significantly from the following literature that focuses on design and convergence of sketched Newton methods. In particular, for many (regression) problems, the objective Hessian can be expressed as $\mathbf{H}=AA^{T}\in\mathbb{R}^{d\times d}$ with a data matrix $A\in\mathbb{R}^{d\times n}$ and $n\geq d$. Then, one can generate a sketch matrix $S\hskip-2.0pt\in\hskip-2.0pt\mathbb{R}^{n\times s}$ and compute the approximate Hessian $\widehat{\mathbf{H}}\hskip-2.0pt=\hskip-2.0ptASS^{T}A^{T}$ilanci2016Iterative developed an iterative Hessian sketch algorithm for solving least-squares problems $\min_{{\bm{x}}}\|A{\bm{x}}-b\|^{2}$ (subject to convex constraints). The authors sketched only the data matrix $A$ rather than both the data matrix $A$ and vector $b$, and established a high-probability convergence result. Lacotte2020Optimal later extended this study by showing the optimal stepsize and convergence rate for Haar sketches. Pilanci2017Newton designed a sketched Newton method that approximates the Hessian using the Johnson–Lindenstrauss transform. Building on this, Agarwal2017Second; Derezinski2019Distributed; Derezinski2020Debiasing; Derezinski2020Precise; Derezinski2021Newton; Lacotte2021Adaptive introduced various sketching methods to explore the trade-off between the computational cost of $\widehat{\mathbf{H}}$ and the convergence rate of the algorithm. In addition to the above series of literature, another type of sketched Newton method is based on Sketch-and-Project framework, where one approximates a generic Hessian inverse $\mathbf{H}^{-1}$ by $S(S^{T}\mathbf{H}S)^{\dagger}S^{T}$ for a sketch matrix $S\in\mathbb{R}^{d\times s}$. See Strohmer2008Randomized; Gower2015Randomized; Luo2016Efficient; Doikov2018Randomized; Gower2019RSN; Derezinski2024Sharp and references therein for the convergence properties of this family of methods.
Compared to deterministic methods for solving Newton systems, such as conjugate gradient and broad preconditioned Krylov (or minimal residual) methods, randomized sketching methods may behave better in terms of improved convergence rates and range of convergence, while requiring less computation and memory (when using suitable sketches) to be scalable and parallelizable (Gower2016Sketch). We refer to Hong2023Constrained for an empirical demonstration of the advantages of sketching solvers over deterministic solvers in the context of SQP methods.
@@ 3 Adaptive Inexact StoSQP Method
$\bullet$ Step 1: Estimate the gradient and Hessian. We realize a sample $\xi_{t}\sim{\mathcal{P}}$ and estimate the gradient $\nabla f_{t}$ and Hessian $\nabla^{2}f_{t}$ of the objective as
@@
We note that the Lagrangian Hessian average in $B_{t}$ is over samples $\{\xi_{0},\ldots,\xi_{t-1}\}$, meaning that the Hessian estimate $\bar{H}_{t}$, which depends on the new sample $\xi_{t}$, will only be used in the $(t+1)$-th iteration. Thus, $B_{t}$ and $\Delta_{t}$ are deterministic given $({\bm{x}}_{t},{\bm{\lambda}}_{t})$. In addition, $\Delta_{t}$ can simply be Levenberg-Marquardt type regularization of the form $\delta_{t}I$ with suitably large $\delta_{t}>0$ (Nocedal2006Numerical). The Hessian regularization is standard for nonlinear problems, together with linear independence constraint qualification (LICQ, Assumption 4.1), ensuring that the quadratic program (3.1) is solvable. For convex problems, we just set $\Delta_{t}={\bm{0}}$, $\forall t$. See Bertsekas1982Constrained; Nocedal2006Numerical for various regularization approaches. Moreover, Na2022Hessian showed that Hessian averaging accelerates the local rate of Newton’s method on unconstrained deterministic problems.
@@
Instead of solving the QP (3.1) exactly, we solve it inexactly by an iterative sketching solver. This approach proves more efficient than deterministic solvers, especially when equipped with suitable sketching matrices (Strohmer2008Randomized; Gower2015Randomized; Pilanci2016Iterative; Pilanci2017Newton; Lacotte2020Optimal). In particular, we generate a random sketching matrix $S\in\mathbb{R}^{(d+m)\times s}$, whose column dimension $s\geq 1$ can also be random, and transform the original large-scale linear system to the sketched, small-scale system as
@@
where $(\cdot)^{\dagger}$ denotes the Moore–Penrose pseudoinverse. One can let $s=1$ (i.e., using sketching vectors) so that $S_{t,j}^{T}K_{t}^{2}S_{t,j}$ reduces to a scalar and the pseudoinverse reduces to the reciprocal.
@@
A significant difference between randomized solvers and deterministic solvers is that the approximation error $\|{\bm{z}}_{t,j}-{\widetilde{\bm{z}}}_{t}\|$ of randomized solvers may not be monotonically decreasing as $j$ increases. This subtlety challenges both inference and convergence analysis. In classical optimization world, it is unanimously agreed that if the search direction is asymptotically close to the exact Newton direction (here ${\widetilde{\bm{z}}}_{t}$), then the algorithm will locally behave just like Newton’s method with a similar convergence rate. A precise characterization is called the Dennis-Moré condition (Dennis1974characterization). In our study, although Lemma LABEL:lem:1 shows that the expected error $\mathbb{E}[\|{\bm{z}}_{t,\tau}-{\widetilde{\bm{z}}}_{t}\|\mid{\bm{x}}_{t},\xi_{t}]$ decays exponentially in $\tau$, ${\bm{z}}_{t,\tau}$ can still be far from ${\widetilde{\bm{z}}}_{t}$ for any $t$ and large $\tau$ in the almost sure sense. In particular, (Patel2021Implicit, Theorem 4.2) proved that only a subsequence of $\{\|{\bm{z}}_{t,\tau}-{\widetilde{\bm{z}}}_{t}\|\}_{\tau=0}^{\infty}$ would decrease monotonically with the subsequence indices being also random.
@@
where $\{\beta_{t},\eta_{t}\}$ are upper and lower bound sequences and $\chi_{t}$ is the adaptivity gap. We do not require specific stepsize selection schemes beyond the condition (3.6) to achieve our online inference goals. The schemes reported in Berahas2021Sequential; Berahas2023Stochastic; Curtis2021Inexact all adhere to the condition (3.6). See (Berahas2021Sequential, Lemma 3.6) and (Curtis2021Inexact, (25, 28)) for details. Their numerical experiments suggest that adaptive random stepsizes offer promising empirical benefits over non-adaptive deterministic stepsizes (i.e., $\chi_{t}=0$). For sake of completeness, we present a selection scheme from Berahas2021Sequential in Appendix LABEL:app:1.